OpenStudy (anonymous):

How to find the derivative of: y= x+ 2xln(2x)

OpenStudy (jdoe0001):

\(\large y= x+ 2xln(2x)\implies \begin{array}{llll} y'=\frac{d}{dx}[&x]+\frac{d}{dx}[&2xln(2x)]\\ &\uparrow &\quad \uparrow \\ &\textit{power rule}&\textit{product rule} \end{array}\)

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

Okay, so I'm confused on the product rule part. Can someone break it down for me?

OpenStudy (anonymous):

The derivative of 2x times ln(2x) + derivative of ln(2x) times 2x

OpenStudy (anonymous):

So that's be 2ln(2x)+ 1/x, right?

OpenStudy (anonymous):

\[2x((1/2x)(2)) + \ln(2x)(2)= 2(\ln(2x)+1)\]

OpenStudy (anonymous):

Thanks!